OpenStudy (anonymous):

Medal + Fan For helping solve this. [9.05] Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64 Please! I will Be fan + Medal This is my last question out of 25 I cannot figure it out!

OpenStudy (amistre64):

you might want to start by showing me what you think the quadratic formula looks like

OpenStudy (amistre64):

there is one part of the formula that is critical to the question, namely: b^2 - 4ac if this setp is negative, there are no "Real" solutions

OpenStudy (amistre64):

*setup

OpenStudy (anonymous):

I don't even know what a quadrant equation is. I really would like to actually learn this stuff. But it's hard!

OpenStudy (espex):

How is it that you are even doing this exercise without knowing what the quadratic equation is?

OpenStudy (anonymous):

I tried to read my lesson, but I did not get it.

OpenStudy (anonymous):

Like I said, I DO want to learn this. I don't just want the answer.

OpenStudy (espex):

You first need to understand that what you are doing is creating an equation that takes the form \[ax^{2}+bx+c=0\]

OpenStudy (espex):

You are familiar with factoring?

OpenStudy (anonymous):

Yes I am familiar with factoring.

OpenStudy (espex):

Okay, give me a quadratic that you can factor.

OpenStudy (anonymous):

x2 − 3x = 28 Is that a quadratic?

OpenStudy (espex):

Almost, you have to be equal to 0 so it would need to be \[x^2-3x-28=0\]

OpenStudy (anonymous):

Okay. Next what?

OpenStudy (espex):

The problem statement wants you to create one that CANNOT be solved by factoring, so what could you do to make this unable to factor?

OpenStudy (espex):

It might help to think about what about it allows you to factor it.

OpenStudy (anonymous):

ax^2 + bx + c = 0?

OpenStudy (anonymous):

This is a form I think. Is it factorable?

OpenStudy (amistre64):

ax^2 + bx + c = 0 is factorable on the Reals if and only if, b^2 -4ac is not a negative value

OpenStudy (amistre64):

the quadratic formula is:\[x=\frac{-b\pm\color{red}{\sqrt{b^2-4ac}}}{2a}\]

OpenStudy (amistre64):

as long as that under the radical is 0 or positive, we can factor it into real numbers

OpenStudy (anonymous):

So using that formula how would we add real numbers in it?

OpenStudy (amistre64):

well, first lets define our favorite values for a and c .... you got any 2 number that make you happy?

OpenStudy (anonymous):

None really, any under 10 will do

OpenStudy (amistre64):

this is yours to work out, you pick ....

OpenStudy (espex):

How about 2 and 7

OpenStudy (espex):

:)

OpenStudy (anonymous):

Okay, my pick... ummmmm. 2 and 4

OpenStudy (amistre64):

lol, great b^2 - 4(2)(4) < 0 b^2 < 4(2)(4) b^2 < 32 you know of any squares less than 32?

OpenStudy (anonymous):

No.

OpenStudy (amistre64):

you know of any squares greater than 32?

OpenStudy (anonymous):

I don't know what you mean?

OpenStudy (amistre64):

b^2 is notation for b squared when we square a number, we multiply it to itself 0*0 = 0 1*1 = 1 2*2 = 4 3*3 = 9 4*4 = 16 5*5 = 25 6*6 = 36 .... any of these squares less than 32 will work

OpenStudy (anonymous):

3*3

OpenStudy (amistre64):

great :) so, a = 2, c = 4, and b= 3 ax^2 + bx + c = 0 2x^2 + 3x + 4 = 0 ... is not factorable under the set of Real numbers

OpenStudy (anonymous):

That makes a little more since!

OpenStudy (amistre64):

good luck :) im just a little curious what "factorable by the quad formula, but not by factoring" might mean.

OpenStudy (espex):

It probably means they want one that you MUST use the quadratic formula for.

OpenStudy (anonymous):

Hey, Thanks so much mate! Very grateful! To the both of you... @eSpeX @amistre64

OpenStudy (amistre64):

nothing NEEDS the quadratic formula tho ... thats just the result of completing the square

OpenStudy (espex):

You're welcome @nicksolis it's nice to find someone who appreciates learning.

OpenStudy (espex):

True that you don't "need" it, but think of the context, you have people learning by factoring and are all of the sudden exposed to the quadratic formula.

OpenStudy (amistre64):

does it mean "integer" factors by chance? or at best rational factors?

OpenStudy (anonymous):

I am not sure.

OpenStudy (amistre64):

if so, then we could simply assert that sqrt(b^2-4ac) not be a perfect square

OpenStudy (amistre64):

\[\frac{-b\pm\sqrt{k^2}}{2a}=\frac{-b-k}{2a},\frac{-b+k}{2a}\] whereas \[\frac{-b\pm\sqrt{p}}{2a}=\frac{-b-\sqrt{p}}{2a},\frac{-b+\sqrt{p}}{2a}~:~p\ne Q\]

OpenStudy (espex):

That would be inline with my thinking. Just something a student could not easily do in their head.

OpenStudy (espex):

Anyway, good luck @nicksolis